Compute the total and annual returns on the described investment.
Five years after buying 150 shares of XYZ stock for $60 per share, you sell the stock for $14,400.
the total return is ___%

Answer by checkley77(12844)

14,400=150*60(1+R)^5
14,400=9,000(1+R)^5
14,400/9,000=(1+R)^5
1.6=(1+R)^5
5THRT(1.6)=1+R
1.09856=1+R
R=1.09856-1
R=.09856 OR 9.856% ANS.
PROOF:
150*60(1+.09856)^5=14,400
9,000(1.09856)^5=14,400
9,000*1.6=14,400
14,400=14,400